Martha brought a $500,000.00 house, paying 10% down, and financing the rest at 5% interest compound monthly for 25 years. Find her monthly payments.
Answer by addingup(3677)

She finances 500,000 - 10% = 450,000
450,000[(0.05/12)(1+(0.05/12))^300]/[1+(0.05/12))^300 - 1] = Payment
450,000[0.00417(3.4848)]/[3.4848 - 1] = Payment
[450,000(0.01453)]/2.4848 = 2,631.40 this is the amount of her monthly payments